OpenStudy (anonymous):

f (26)x = 1, what is the value of x? Explain your answer. (

OpenStudy (anonymous):

please help!!!!!!!!

OpenStudy (anonymous):

\((2^6)^x = 1\), Remember : Anything (except \(0\)) raised to the power \(0\) is \(1\)..

OpenStudy (anonymous):

So: \[\large (2^6)^x = 1 \\ (2^6)^x = (2^6)^0 \implies \color{green}{x = 0}\]

OpenStudy (anonymous):

Using: \[(a)^x = (a)^y \implies x = y\]

OpenStudy (anonymous):

so to get 1 the missing exponent is 0?

OpenStudy (anonymous):

@eta

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

\[(2^6)^0 = 1\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

can you help me with a few more please? @eta